the square of katie's age six year ago is equal to her age in six years time. find her present age
Answer by Maths68(1474)

Let
Present age of Katie = k years old

Six years ago
Katie was = k-6 years old
In Six years
Katie will be = k+6 years old
The square of katie's age six year ago is equal to her age in six years time
(k-6)^2=k+6
k^2-12k+36=k+6
k^2-13k+30=0
Solve for K
k^2-10k-3k+30=0
k(k-10)-3(k-10)=0
(k-3)(k-10)=0
k-3=0 or k-10=0
k=3 (unacceptable) or k=10
k=10

Present age of Katie = k = 10 years old